a bus is moving at 8m/s when passenger walks forward at 1.5m/s relative to the bus . how fast is this passenger moving as seen and measured by (a) a person waiting at the bus stop and (b) the driver of the bus?


1
Expert's answer
2022-06-07T09:24:46-0400

Given:

"v_b=8.0\:\rm m/s"

"v_p=1.5\:\rm m/s"

a)


"V_r=v_b+v_p=8.0+1.5=9.5\:\rm m/s"



(b)


"V_r=v_p=1.5\:\rm m/s"